OpenStudy (anonymous):

someone help me please :c Solve the following quadratic equation: (x + 3)^2 = 16. A. x = -3, -4 B. x = 1 C. x = -1, 7 D. x = 1, -7

hero (hero):

(x + 3)^2 = 16 x + 3 = ± 4 x = ± 4 - 3 x = -7 x = 1

OpenStudy (anonymous):

thank u!

hero (hero):

Do you understand the steps?

OpenStudy (anonymous):

not at all. im trying to understand bt it just doesn't click...

hero (hero):

You start with (x + 3)^2 = 16. Then you take the square root of both sides: \[\sqrt{(x + 3)^2} = \sqrt{16}\] Remember that square root and square are inverses of each other so they cancel \[x + 3 = \pm 4\] The plus or minus sign results from the fact that the square root of 16 is 4 and -4 because \((-4)^2 = 16\) and \(4^2 = 16\) So now you split \(x + 3 = ±4\) into two equations: \(x + 3 = 4\) \(x + 3 = -4\) From there you solve for x